OpenStudy (anonymous):

solve for a. 5a + 7 =b

OpenStudy (saifoo.khan):

Step 1; Shift 7 on the other side. Step 2: Shift 5 on the other side.

OpenStudy (anonymous):

ok like 5a + 7 = b to 7a + 5 = b

OpenStudy (hba):

@kittylove Follow me 5a + 7 =b 1)subtract 7 by both sides 2) divide 5 by both sides

OpenStudy (anonymous):

5-7 = -2 right

OpenStudy (hba):

No See, 5a + 7-7 =b-7 Now simplify this and show me